For a system of mass m=0.05kg and k=50Nm-¹ at time t=0, the mass is observed to be displaced 60mm to the right of its equilibrium position and to be moving to the right at the speed of 2.5ms-¹. Calculate:

a)the amplitude

b)the phase constant

c)the energy


1
Expert's answer
2021-09-02T10:42:48-0400

a)


"kA^2=kx^2+mv^2\\50A^2=50(0.06)^2+(0.05)(2.5)^2\\A=0.099\ m"

b)


"0.06=0.099\cos {\phi}\\\phi= 0.92\ rad"

c)


"E=0.5(50)(0.099)^2=0.245\ J"